A volume 8m3 of an ideal gas is held at a constant temperature and pressure 1atm. If the volume changes to 25 m3, the pressure will be ____ atm.
Answer #1

ideal gas equation

PV=nRT

P is pressure

V is volume

n is number of moles

R is gas constant

T is temperature in kelvin

since T is constant nRT term is constant as number of moles(n) will remain same ,R is itself constant,T is said to be constant

so P1V1=P2V2

8*1=P2*25

P2=8/25=0.32 atm